OpenStudy (anonymous):

square root 2x+6 - square root x+4 =1

OpenStudy (anonymous):

\[\sqrt{2x+6}-\sqrt{x+4}=1\]

OpenStudy (anonymous):

?

OpenStudy (ajprincess):

\[\sqrt{2x+6}-\sqrt{x+4}=1\] Square both sides. \[(\sqrt{2x+6}-\sqrt{x+4})^2=1^2\] \(2x+6-2\sqrt{2x+6}\sqrt{x+4}+x+4=1\) \(((a-b)^2=a^2-2ab+b^2)\) \(3x+10-1=2\sqrt{2x+6}\sqrt{x+4}\) \(3x+9=\sqrt{(2x+6)(x+4)}\) Square both sides. \((3x+9)^2=(\sqrt{(2x+6)(x+4)})^2\) \(9x^2+54x+81=2x^2+14x+24\) \(9x^2+54x+81-2x^2-14x-24=0\) \(7x^2+40x+57=0\) Nw try to solve this quadratic equation. Does that help? @Lii

OpenStudy (anonymous):

thanks alot, it helped alot to get started

OpenStudy (ajprincess):

Welcome:) To solve the quadratic equation, u can use quadratic formula. \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] the values of a.b and c can be found by comparing the equation \(7x^2+40x+57=0\) with the general equation \(ax^2+bx+c=0\)

OpenStudy (anonymous):

i got -2.7143 and -3

OpenStudy (ajprincess):

perfect:)